OpenStudy (anonymous):

4. The legs of a right triangle measure 24 feet and 45 feet. What is the length of the hypotenuse? (1 point)51 ft 69 ft 38.07 ft 8.31 ft 5. Which of these cannot represent the lengths of the sides of a right triangle? (1 point)3 ft, 4 ft, 5 ft 6 in., 8 in., 10 in. 16 cm, 63 cm, 65 cm 8 m, 9 m, 10 m

OpenStudy (anonymous):

can some one please answer these questions I need them today!!!!!!!!!!!!!!!!

OpenStudy (lexipoo1998):

The first one is 51. hold on and il will do the other two

OpenStudy (lexipoo1998):

*other one

OpenStudy (anonymous):

k thanks :)

OpenStudy (lexipoo1998):

16,63,65 cannot be the lengths of a triangle. Do you know why?

OpenStudy (anonymous):

have no clue......

OpenStudy (lexipoo1998):

Would you like me to explain it a nic and simple way?

OpenStudy (anonymous):

yess thank and thank you so much

OpenStudy (lexipoo1998):

No problem and ok. The way i remeber is that the two of the three numbers given cannot equal less or more then them added up or subtracted form eachother. It sounds complicated but it is really not. For example: I'll use your problem 3ft,4ft,5ft 3+4 = 7 so a leg can not be more than 7 which in this case it is not 5-3 = 2 so a leg can not be less than 2 which it is also not so it is a triangle BUT with the one that is incorrect you would do 63-16 = 47 so without even doing the adding part you know that the leg 16 can not be since once you do the subtraction you see in order to be a triangle the leg can not be less that 47ft Do you kinda get it know?

OpenStudy (anonymous):

yess again thank you so much buttt I got 2 go now my teachers coming in like right now bye.:)

OpenStudy (lexipoo1998):

Lol ok good luck with your future math stuff. BYE